Consider a vector F=4i-3j. another vector that is perpendicular to F is-
 
can you show me the procedure to do this question?

Grade:12

8 Answers

Trina
8 years ago
the dot product must be 0
let the perpendicular force be p = xi+yj
(xi+yj).(4i-3j)=0
4x-3y=0
x=3/4y
p=3/4i+j or p=3i+4j
Kirandeep
6 years ago
K cap is answer of the abovery question. It will be a unit vector along z -axis as F vector lies in X-Y plane.
Jashan
6 years ago
Consider a vector F =4icap-3jcap.the vector perpendicular to vector F is??What is the unit vector of F
Sangeetha
5 years ago
The dot product must be 0  
Let the perpendicular force be xi+yj 
(4i-3j) . (Xi+yj)=0 
4x-3y=0 
4x=3y 
P=3i +4j 
i  j    k 
3 4  0 
4 -3  0 
By using cross product formula we gets 7k^
mahima
5 years ago
in cartesian form of vectors ,all vectors are perpendicular to each other  .
we have given,
f=4i-3j
according to cartesian form,k will be perpendicular to it
i.e.4-(-3)=7k
 
Aakansh raj
4 years ago
The ans is 7k cap.
i know phy is tough but regular practice will make it easy
the more question u solve more knowledge u will get
Ajeet Tiwari
3 years ago
hello student

the dot product must be 0

let the perpendicular force be p = xi+yj

(xi+yj).(4i-3j)=0

4x-3y=0

x=3/4y

p=3/4i+j or p=3i+4j

Hope it helps
Thankyou